# Stronger Together: On-Policy Reinforcement Learning for Collaborative LLMs

### v2

@@ Abstract
Multi-Agent System (MAS) and Reinforcement Learning (RL) are both widely adopted to improve large language model (LLM) agentic performance. MAS strengthens task-specialized performance via role-based orchestration; RL leverages environment rewards to train stronger policies, such as Group Relative Policy Optimization (GRPO)-style optimization. Yet applying on-policy RL training to MAS is underexplored. While promising, it poses several challenges. On the algorithm side, Standard GRPO grouping assumptions fail in MAS because prompts differ by role and turn. On the system side, the training system needs to support MAS-workflow-based rollouts and on-policy updates for both single and multiple policy models. To address these issues, we introduce AT-GRPO, consisting of (i) an Agent- and Turn-wise grouped RL algorithm tailored for MAS and (ii) a system to support both single-policy and multi-policy training. Across game, plan, coding, and math tasks, AT-GRPO  demonstrates substantial performance gains across diverse domains. Especially on long-horizon planning tasks, AT-GRPO boosts accuracy from a 14.0–47.0% single-agent RL baseline to 96.0–99.5%. Furthermore, it improves reasoning performance, with an average gain of 3.87–7.62% on coding and 9.0-17.93% on math.  [^1] Code and environments are available at: https://github.com/pettingllms-ai/PettingLLMs
@@ 1 Introduction
Large Language Model (LLM) agents are task-specific workflows (Yao et al., 2023; Xi et al., 2023; Wang et al., 2023b) that utilize LLMs as key components for decision making (Shinn et al., 2023), action taking (Wang et al., 2023a), and tool use (Qian et al., 2025; Schick et al., 2023). LLM agents have demonstrated strong promises across various application domains, such as embodied control (Ahn et al., 2022; Wang et al., 2023a), software engineering (Tao et al., 2024; Yu et al., 2025), expert drug discovery (Liu et al., 2024; Inoue et al., 2024), and scientific ideation and hypothesis testing (Ghafarollahi and Buehler, 2024).
Today, two complementary approaches are widely used to improve the performance of LLM agents: multi-agent systems (MAS) and reinforcement learning (RL). RL treats the LLM as a policy and iteratively updates its weights to strengthen decision-making: at each iteration, the current model interacts with the environment, collects rule-based rewards, and then computes a policy optimization loss to update the parameters (Shao et al., 2024). In practice, this workflow requires a training stack that supports both scalable rollouts and online updates, e.g., VERL (Sheng et al., 2025) and AReaL (Fu et al., 2025). MAS typically employs prompt-only augmentation on a shared LLM policy for role-based coordination; practical deployments instantiate diverse workflows. Recent studies (Belcak et al., 2025; Chen et al., 2024; Wang et al., 2024) further highlight the potential benefits of role-specialized MAS, which adopts distinct models for different roles, enabling role-specialized policies in inference. However, the effectiveness of RL training on role-specialized MAS is underexplored.
@@
A natural next step is to combine the two: using RL to train MAS, such that we gain both stronger learned policies, role-specialized collaboration. However, bringing RL into MAS raises two coupled challenges. First, training a MAS may require concurrently launching multiple models, orchestrating inter-agent environment interactions, and performing independent on-policy parameter updates. But most existing on-policy RL frameworks for LLM agents only support a single model (Volcano Engine, 2025; Sheng et al., 2024; Fu et al., 2025). Second, rollouts from MAS are difficult to group. The advantage must be conditioned on interaction history and role to ensure fair credit assignment. Group-based RL objectives designed for a single agent (Volcano Engine, 2025; Qian et al., 2025; Feng et al., 2025) are not directly applicable to MAS.
@@ 2 Related Work
RL for LLM Agentic Training. RL has become a key technique for LLMs agent training, using group-relative and rule-based rewards to enhance reasoning, long-horizon planning, game, and tool use (Shao et al., 2024; Wang et al., 2025b; Qian et al., 2025; Hu et al., 2025). These approaches, however, predominantly operate within a single-agent framework. Although effective for certain benchmarks, this paradigm offers limited potential for improvement, as it relies on a single agent for planning and neglects the inherent advantages of MAS for complex coordination and specialization, thereby constraining further breakthroughs.
Role-sharing vs. Role-specialized Policies in MAS. A predominant approach in LLM-based MAS centers on a role-sharing architecture, where a single policy is shared across all agents. In these frameworks, such as AutoGen (Wu et al., 2023) and MetaGPT (Hong et al., 2024), role-specific behavior is elicited at inference time via prompt augmentation. More recently, research has begun to explore role-specialized policies. This shift is motivated by the observation that a single LLM’s performance exhibits significant variance across domains (Chen et al., 2024; Wang et al., 2024; Belcak et al., 2025). Consequently, assigning distinct and more suitable models to specialized roles, as demonstrated by Ye et al. (2025); Belcak et al. (2025), has emerged as a promising direction for enhancing performance. Despite this architectural evolution, recent surveys (Cemri et al., 2025; Guo et al., 2024) indicate that most studies focus on inference-time design, leaving the potential of training MAS policies with RL largely underexplored.
RL Training for MAS. A growing body of work tries to bring RL to MAS, but most efforts are confined to a single interaction and role-sharing policy pattern. MAPoRL (Park et al., 2025a; b), CoRY Ma et al. (2024) train LLMs as each agent proposes and revises answers to the same query within a shared discussion and debate with each other, CURE (Wang et al., 2025a) co-evolves a Coder and a Unit-Tester with a role-sharing policy for code generation; SPIRAL (Liu et al., 2025) trains via self-play on zero-sum games using a single LLM; and MHGPO (Chen et al., 2025a) targets retrieval-augmented generation, coupling group-based objectives with a role-sharing policy around retrieval, routing, and response selection. Compared with these approaches, our study is comprehensive: we evaluate across diverse MAS workflows from different domains, and comprehensively compare role-sharing versus role-specific policies.
@@ 3 Preliminaries
Figure 2: MAS workflow across different domains. (a) Role-based coordination: code generation via a coder–tester loop. (b) Different task-specific workflows for Game/Plan, Code, and Math; see Sec. 5.1 and Appendix A.2.2 for workflow details.
@@
We model an $N$-agent, LLM-based multi-agent system as a Markov game $\mathcal{M}=(\mathcal{S},\{\mathcal{A}_{i}\}_{i=1}^{N},\mathcal{T},\{r_{i}\}_{i=1}^{N},T,H)$, where $\mathcal{S}$ is the state space; $\mathcal{A}_{i}$ is the action space of agent $i$; $\mathcal{T}$ is a transition map $s_{t+1}=\mathcal{T}(s_{t},a_{1,t},\ldots,a_{N,t})$; The reward for agent $i$ is given by $r_{i}:\mathcal{A}_{i}\to[0,1]$, and the turn horizon $T$, the optimization step horizon $H$. At each turn $t$, agent $i$ receives an observation summarizing the environment state and interaction history $h_{t}$, $o_{i,t}\!=\!o_{i}(s_{t},h_{t})$. Each agent $i$ is implemented with a role-specific prompt template $\mathsf{P}_{i}(\cdot)$. Let $\Theta=\{\theta^{(m)}\}_{m=1}^{M}$ denote the set of LLM parameter vectors, with $1\leq M\leq N$, and let $\sigma:\{1,\ldots,N\}\!\to\!\{1,\ldots,M\}$ assign each agent to an LLM. We treat one LLM rollout (a token sequence) as a single macro-action $a_{i,t}$. A turn is one full interaction in which all agents emit macro-actions to the environment. A step denotes one optimization update to the parameter set $\Theta$ during training.
@@
Following prior work (Wang et al., 2025a; Ahn et al., 2022; Chen et al., 2025b), we employ domain-specific MAS workflows, as shown in Fig. 2. Our experiments confirm that this prompt-only method outperforms a single-agent baseline (see Tab. 2 and 2 in Sec. 5.2).
@@
Methods for LLM agentic training with group-relative advantages (Feng et al., 2025; Wang et al., 2025b; Qian et al., 2025) operate by first sampling $K$ candidate actions $\{a_{t}^{(k)}\}_{k=1}^{K}$ for a given prompt. Each action is evaluated to obtain a rule-based reward $R(a_{t}^{(k)})$, forming a comparison group: $G=\big\{\,(a_{t}^{(1)},R(a_{t}^{(1)})),\;\ldots,\;(a_{t}^{(K)},R(a_{t}^{(K)}))\,\big\}.$ For each action $a_{t}^{(k)}$ in this group, the relative advantage is then defined as its mean-centered and normalized return.
$A^{E}\!\big(a_{t}^{(k)}\big)=\frac{\,R(a_{t}^{(k)})-\mathrm{mean}\!\left(\{\,R(a_{t}^{(\ell)})\,\}_{\ell=1}^{K}\right)\,}{\,F_{\mathrm{norm}}\!\left(\{\,R(a_{t}^{(\ell)})\,\}_{\ell=1}^{K}\right)\,}$, (1)
where $F_{\mathrm{norm}}(\cdot)$ is the sample standard deviation.
@@
We distinguish between two optimization regimes, role-sharing and role-specialized, both of which initialize policies from the same base model. During rollouts, each agent $i$ generates a dataset $\mathcal{D}_{i}$, which consists of sample groups. A single group $g$ is composed of a shared observation context $o_{g}$ and $K$ candidate actions with their corresponding advantages, denoted as $g=\{i,a_{g}^{(c)},A_{g}^{(c)}\}_{c=1}^{K}$. The core difference between the two regimes lies in how the training data is batched. A minibatch $\mathcal{B}_{m}$ for a specific policy $\theta^{(m)}$ is constructed by pooling the datasets from all agents assigned to it:
$\mathcal{B}_{m}=\bigcup_{i\,:\,\sigma(i)=m}\mathcal{D}_{i}.$ (2)
Each policy is then updated using a GRPO-style objective over its corresponding minibatch:
$\mathcal{L}(\theta^{(m)})\;=\;-\,\mathbb{E}_{g\in\mathcal{B}_{m}}\!\left[\frac{1}{K}\sum_{c=1}^{K}\log\pi_{\theta^{(m)}}\!\big(a_{g}^{(c)}\mid\mathsf{P}_{i}(o_{g})\big)\;A_{g}^{(c)}\right].$ (3)
Role-sharing policy (M=1M{=}1): All agents share a single policy $\theta^{1}$. The training batch is the union of data from all agents, $\mathcal{B}_{1}=\bigcup_{i=1}^{N}\mathcal{D}_{i}$, and is used for a single joint update: $\theta^{1}\leftarrow\theta^{1}-\eta\nabla_{\theta^{1}}\mathcal{L}(\theta^{1})$.
@@ 4.1 Algorithm Design: AT-GRPO
GRPO’s advantage calculation (Eq. 1) hinges on a fair comparison among all candidates within a group. This fairness is enforced by the reward mechanism itself. As illustrated in Fig. 3 (top), token-level scoring assigns credit to the generated response tokens (Reward Mask=1), while the prompt tokens receive no credit (Reward Mask=0). Since the advantage is determined solely by the quality of the response, a valid and fair comparison is only possible when all responses in a group originate from an identical prompt. Consequently, single-agent LLM-RL methods(Wang et al., 2025b; Qian et al., 2025; Feng et al., 2025) typically form groups by sampling multiple responses to the same question.
In MAS, however, a “prompt” is not only a question description, but also embeds the role-specific context and cross-agent interaction history. For example, in code debugging (Fig. 3, middle), the turn-2 refinement prompt already contains the turn-1 code, unit tests, and role-specific prompt format, so prompts differ across turns and roles. Thus, grouping by “same question” no longer yields comparable samples. We therefore adopt agent-wise and turn-wise grouping: candidates share the same role and turn position, ensuring prompt identity for valid GRPO advantage comparisons.
@@
7:   Get initial state $s_{0,e}$.
8:   for $t=0$ to T1T-1 do
@@
10:      $\forall c\in\{1,\dots,K\}$, sample $a^{(c)}_{i,t,e}\sim\pi_{\theta^{(\sigma(i))}}(\cdot\mid o_{i,t,e};\,T_{\text{samp}})$; compute $r^{(c)}_{i,t,e}$(Eq. 4)
11:      Define group key $g\leftarrow\text{hash}(e,i,t)$ and compute advantages $\{A^{(c)}_{i,t,e}\}_{c=1}^{K}$ (Eq. 1). (Agent- and turn-wise grouping.)
12:      Append $(g,o_{i,t,e},\{a^{(c)}_{i,t,e}\}_{c=1}^{K},\{A^{(c)}_{i,t,e}\}_{c=1}^{K})$ to $\mathcal{D}_{i}$.
13:      $c^{\star}\leftarrow\arg\max_{c}r^{(c)}_{i,t,e}$; $a_{i,t,e}\leftarrow a^{(c^{\star})}_{i,t,e}$. (Tree-structured sampling.)
14:     end for
15:     $s_{t+1,e}\leftarrow\mathcal{T}\big(s_{t,e},a_{1,t,e},\dots,a_{N,t,e}\big)$.
16:     if $\mathcal{I}_{\mathrm{term}}(s_{t+1},e)$ then break
17:     end if
18:   end for
@@
20:/*– Phase 2: Per-Model Policy Update –*/
21:  for each model $m\in\{1,\dots,M\}$ in parallel do
22:   Construct per-model batch $\mathcal{B}_{m}$ using Eq. 2.
23:   Compute loss $\mathcal{L}(\theta^{(m)})$ on $\mathcal{B}_{m}$ using Eq. 3 and update policy $m$.
@@
Inspired by mixed-reward designs in cooperative Multi-Agent RL (Mao et al., 2020; Sheikh and Bölöni, 2020), we assign credit using a mixture of global and local rewards. At each turn $t$, the environment provides a global team reward $r^{\mathrm{team}}$ and an agent-specific local reward $r^{\mathrm{loc}}_{i}$ that evaluates its subtask performance. These components are combined using a hyperparameter $\alpha$ to form the final reward for agent $i$:
$r_{i,t}\;=\;\alpha\,r^{\mathrm{team}}\;+\;r^{\mathrm{loc}}_{i}$ (4)
This formulation balances a shared team objective with role-specific incentives. For instance, in a coder-tester MAS, the team reward $r^{\mathrm{team}}$ is the pass rate of the generated program on a set of golden unit tests. The local rewards $r^{\mathrm{loc}}_{i}$ are tailored to each role: the coder is rewarded for its own code’s pass rate, while the tester is rewarded based on the pass rate of a golden reference implementation against its generated tests. Detailed reward designs for all tasks are provided in Appendix A.2.1.

### v4

# Stronger-MAS: Multi-Agent Reinforcement Learning for Collaborative LLMs

@@ Abstract
Multi-Agent System (MAS) and Reinforcement Learning (RL) are both widely adopted to improve large language model (LLM) agentic performance. MAS strengthens task-specialized performance via role-based orchestration; RL leverages environment rewards to train stronger policies, such as Group Relative Policy Optimization (GRPO)-style optimization. Yet applying on-policy RL training to MAS is underexplored. While promising, it poses several challenges. On the algorithm side, Standard GRPO grouping assumptions fail in MAS because prompts differ by role and turn. On the system side, the training system needs to support MAS-workflow-based rollouts and on-policy updates for both single and multiple policy models. To address these issues, we introduce AT-GRPO, consisting of (i) an Agent- and Turn-wise grouped RL algorithm tailored for MAS and (ii) a system to support both single-policy and multi-policy training. Across game, plan, coding, and math tasks, AT-GRPO  demonstrates substantial performance gains across diverse domains. Especially on long-horizon planning tasks, AT-GRPO boosts accuracy from a 14.0–47.0% single-agent RL baseline to 96.0–99.5%. Furthermore, it improves reasoning performance, with an average gain of 3.87–7.62% on coding and 9.0-17.93% on math.
@@ 1 Introduction
Large Language Model (LLM) agents are task-specific workflows (Yao2023ReAct; xi2023rise; Wang2023LLMAgentsSurvey) that utilize LLMs as key components for decision making (shinn2023reflexion), action taking (wang2023voyager), and tool use (Qian2025ToolRL; schick2023toolformer). LLM agents have demonstrated strong promises across various application domains, such as embodied control (ahn2022saycan; wang2023voyager), software engineering (tao2024magis; yu2025orcaloca), expert drug discovery (liu2024drugagent; inoue2024drugagent), and scientific ideation and hypothesis testing (ghafarollahi2024sciagents).
Today, two complementary approaches are widely used to improve the performance of LLM agents: multi-agent systems (MAS) and reinforcement learning (RL). RL treats the LLM as a policy and iteratively updates its weights to strengthen decision-making: at each iteration, the current model interacts with the environment, collects rule-based rewards, and then computes a policy optimization loss to update the parameters (shao2024deepseekmath). In practice, this workflow requires a training stack that supports both scalable rollouts and online updates, e.g., VERL (sheng2025hybridflow) and AReaL (fu2025areal). MAS typically employs prompt-only augmentation on a shared LLM policy for role-based coordination; practical deployments instantiate diverse workflows. Recent studies (belcak2025slm_agentic; chen2024ioa; wang2024moa) further highlight the potential benefits of role-specialized MAS, which adopts distinct models for different roles, enabling role-specialized policies in inference. However, the effectiveness of RL training on role-specialized MAS is underexplored.
@@
A natural next step is to combine the two: using RL to train MAS, such that we gain both stronger learned policies, role-specialized collaboration. However, bringing RL into MAS raises two coupled challenges. First, training a MAS may require concurrently launching multiple models, orchestrating inter-agent environment interactions, and performing independent on-policy parameter updates. But most existing on-policy RL frameworks for LLM agents only support a single model (volcengine2025verl; sheng2024hybridflow; fu2025areal). Second, rollouts from MAS are difficult to group. The advantage must be conditioned on interaction history and role to ensure fair credit assignment. Group-based RL objectives designed for a single agent (volcengine2025verl; Qian2025ToolRL; feng2025group) are not directly applicable to MAS.
@@ 2 Related Work
Role-sharing vs. Role-specialized Policies in MAS. A predominant approach in LLM-based MAS centers on a role-sharing architecture, where a single policy is shared across all agents. In these frameworks, such as AutoGen (wu2023autogen) and MetaGPT (hong2024metagpt), role-specific behavior is elicited at inference time via prompt augmentation. More recently, research has begun to explore role-specialized policies. This shift is motivated by the observation that a single LLM’s performance exhibits significant variance across domains (chen2024ioa; wang2024moa; belcak2025slm_agentic). Consequently, assigning distinct and more suitable models to specialized roles, as demonstrated by ye2025xmas; belcak2025slm_agentic, has emerged as a promising direction for enhancing performance. Despite this architectural evolution, recent surveys (pan2025whydomasfail; guo2024llmma) indicate that most studies focus on inference-time design, leaving the potential of training MAS policies with RL largely underexplored.
RL Training for MAS. RL has become a key technique for LLMs agent training, using group-relative and rule-based rewards to enhance reasoning, long-horizon planning, game, and tool use (feng2025group; wang2025ragen; Qian2025ToolRL; hu2025lmgame). These approaches, however, predominantly operate within a single-agent framework. While a growing body of work attempts to extend RL to Multi-Agent Systems (MAS), most efforts remain confined to limited interaction settings or fixed role structures. For instance, CURE (wang2025cure) focuses on co-evolving a Coder and Unit-Tester using a role-sharing policy specifically for code generation. Similarly, SPIRAL (liu2025spiral) employs self-play in zero-sum games using a single LLM, while MHGPO (chen2025mhgpo) targets retrieval-augmented generation. MAPoRL (park2025maporl; park2025maporl2) and CoRY ma2024cory train LLMs within fixed, homogeneous-role debate workflows. More recent works also exhibit limitations: MARFT Liao2025MARFT  [^1] We compare against MARFT v3, the latest preprint available prior to the completion of this work. restricts agents to single-turn sequential interactions, and MARTI zhang2025marti merely introduces basic single-agent RL algorithms (e.g., GRPO) to the MAS setting. Critically, the implemention for these works is concentrated on the single domain of math. We include a more comprehensive comparison of these related works in Appendix A. In contrast, our study offers a more comprehensive solution. We propose a general algorithm designed for multi-turn, multi-agent environments. Unlike prior works, we conduct a thorough analysis and evaluation of both shared and role-specific policies across diverse MAS workflows and varying domains.
@@ 3 Preliminaries
Figure 2: MAS workflow across different domains. (a) Role-based coordination: code generation via a coder–tester loop. (b) Different task-specific workflows for Game/Plan, Code, and Math; see Sec. 5.1 and Appendix C.2 for workflow details.
@@
The $N$-agent LLM system is modeled as a Markov game $\mathcal{M}=(\mathcal{S},\{\mathcal{A}_{i}\}_{i=1}^{N},\mathcal{T},\{r_{i}\}_{i=1}^{N},T,H)$, where $\mathcal{S}$ is the state space; $\mathcal{A}_{i}$ is the action space of agent $i$; The transition function $\mathcal{T}$ induces intra-turn micro-transitions where $s_{t,0}=s_{t}$ and $s_{t,i}=\mathcal{T}(s_{t,i-1},a_{t,i},i)$, culminating in $s_{t+1}=s_{t,N}$. The reward for agent $i$ is given by $r_{i}:\mathcal{A}_{i}\to[0,1]$, and the turn horizon $T$, the optimization step horizon $H$. At each turn $t$, agent $i$ receives an observation summarizing the environment state and interaction history $h_{t}$, $o_{t,i}\!=\!o_{i}(s_{t},h_{t})$. Each agent $i$ is implemented with a role-specific prompt template $\mathsf{P}_{i}(\cdot)$. Let $\Theta=\{\theta^{(m)}\}_{m=1}^{M}$ denote the set of LLM parameter vectors, with $1\leq M\leq N$, and let $\sigma:\{1,\ldots,N\}\!\to\!\{1,\ldots,M\}$ assign each agent to an LLM. We treat one LLM rollout (a token sequence) as a single macro-action $a_{t,i}$. A turn is one full interaction in which all agents emit macro-actions to the environment. A step denotes one optimization update to the parameter set $\Theta$ during training.
@@
Following prior work (wang2025cure; ahn2022saycan; Chen2025CodeSteer), we employ domain-specific MAS workflows, as shown in Fig. 2. Our experiments confirm that this prompt-only method outperforms a single-agent baseline (see Tab. 2 and 2 in Sec. 5.2).
@@
Methods for LLM agentic training with group-relative advantages (feng2025group; wang2025ragen; Qian2025ToolRL) operate by first sampling $K$ candidate actions $\{a_{t}^{(c)}\}_{c=1}^{K}$ for a given prompt. Each action is evaluated to obtain a rule-based reward $R(a_{t}^{(c)})$, forming a comparison group: $G=\big\{\,(a_{t}^{(1)},R(a_{t}^{(1)})),\;\ldots,\;(a_{t}^{(K)},R(a_{t}^{(K)}))\,\big\}.$ For each action $a_{t}^{(c)}$ in this group, the relative advantage is then defined as its mean-centered and normalized return.
$A_{g}\!\big(a_{t}^{(c)}\big)=\frac{\,R(a_{t}^{(c)})-\mathrm{mean}\!\left(\{\,R(a_{t}^{(c)})\,\}_{c=1}^{K}\right)\,}{\,F_{\mathrm{norm}}\!\left(\{\,R(a_{t}^{(c)})\,\}_{c=1}^{K}\right)\,}$, (1)
@@
We distinguish between two optimization regimes, role-sharing and role-specialized, both of which initialize policies from the same base model. During rollouts, each agent $i$ generates a dataset $\mathcal{D}_{i}$, which consists of sample groups. A single group $g$ is composed of a shared observation context $o_{g}$ and $K$ candidate actions with their corresponding advantages, denoted as $g=\{i,a_{g}^{(c)},A_{g}^{(c)}\}_{c=1}^{K}$. The core difference between the two regimes lies in how the training data is batched. A minibatch $\mathcal{B}_{m}=\bigcup_{i\,:\,\sigma(i)=m}\mathcal{D}_{i}.$ for a specific policy $\theta^{(m)}$ is constructed by pooling the datasets from all agents assigned to it:
$\mathcal{L}(\theta^{(m)})\;=\;-\,\mathbb{E}_{g\in\mathcal{B}_{m}}\!\left[\frac{1}{K}\sum_{c=1}^{K}\min\!\Big(r_{g}^{(c,m)}(\theta^{(m)})\,A_{g}^{(c)}\,,\operatorname{clip}\!\big(r_{g}^{(c,m)}(\theta^{(m)}),\,1-\varepsilon,\,1+\varepsilon\big)\,A_{g}^{(c)}\Big)\right]$ (2)
where $r(\theta)=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}$. Role-sharing policy (M=1M{=}1): All agents share a single policy $\theta^{1}$. The training batch is the union of data from all agents, $\mathcal{B}_{1}=\bigcup_{i=1}^{N}\mathcal{D}_{i}$, and is used for a single joint update: $\theta^{1}\leftarrow\theta^{1}-\eta\nabla_{\theta^{1}}\mathcal{L}(\theta^{1})$.
@@ 4.1 Algorithm Design: AT-GRPO
GRPO’s advantage calculation (Eq. 1) hinges on a fair comparison among all candidates within a group. This fairness is enforced by the reward mechanism itself. As illustrated in Fig. 2 (top), token-level scoring assigns credit to the generated response tokens (Reward Mask=1), while the prompt tokens receive no credit (Reward Mask=0). Since the advantage is determined solely by the quality of the response, a valid and fair comparison is only possible when all responses in a group originate from an identical prompt. Consequently, single-agent LLM-RL methods(wang2025ragen; Qian2025ToolRL; feng2025group) typically form groups by sampling multiple responses to the same question.
In MAS, however, a “prompt” is not only a question description, but also embeds the role-specific context and cross-agent interaction history. For example, in the code tasks depicted in Fig. 2 (a) , where the workflow entails a coder-tester loop: one agent synthesizing code and the other creating unit tests, and they iteratively refine the output until alignment (Fig. 3, top), the turn-2 refinement prompt already contains the turn-1 code, unit tests, and role-specific prompt format, so prompts differ across turns and roles. We therefore adopt agent-wise and turn-wise grouping as a natural extension of tabular-wise group-normalized advantages in GiGPO (feng2025group) to the multi-agent setting: candidates share the same role and turn position, ensuring prompt identity for valid GRPO advantage comparisons.
@@
7:   for $t=0$ to T1T-1 do
8:     $s_{t,0,e}\leftarrow s_{t,e}$ $\triangleright$ Initialize micro-step state
@@
10:      $\forall c\in\{1,\dots,K\}$, $a^{(c)}_{t,i,e}\sim\pi_{\theta^{(\sigma(i))}}(\cdot\mid o_{t,i,e};\,T_{\text{samp}})$; compute $r^{(c)}_{t,i,e}$ (Eq. 3)
11:      Define group key $g\leftarrow\text{hash}(e,i,t)$ and compute advantages $\{A^{(c)}_{g}\}_{c=1}^{K}$ (Eq. 1).
12:      Append $(g,o_{t,i,e},\{a^{(c)}_{t,i,e}\}_{c=1}^{K},\{A^{(c)}_{g}\}_{c=1}^{K})$ to $\mathcal{D}_{i}$.
13:      $c^{\star}\leftarrow\arg\max_{c}r^{(c)}_{t,i,e}$; $a_{t,i,e}\leftarrow a^{(c^{\star})}_{t,i,e}$. (Tree-structured sampling.)
14:      $s_{t,i,e}\leftarrow\mathcal{T}\big(s_{t,i-1,e},a_{t,i,e},i\big)$ $\triangleright$ Agent-wise micro-transition
15:     end for
16:     $s_{t+1,e}\leftarrow s_{t,N,e}$ $\triangleright$ End-of-turn state
17:     if $\mathcal{I}_{\mathrm{term}}(s_{t+1,e})$ then break
18:     end if
@@
20:  end for
21:/*– Phase 2: Per-Model Policy Update –*/
22:  for each model $m\in\{1,\dots,M\}$ in parallel do
23:   Construct per-model batch $\mathcal{B}_{m}$, loss $\mathcal{L}(\theta^{(m)})$ on $\mathcal{B}_{m}$ using Eq. 2 and update policy $m$.
@@
Inspired by mixed-reward designs in cooperative Multi-Agent RL (Mao2020RewardDesignMARL; Sheikh2020DEMADDPG), we assign credit using a mixture of global and local rewards. At each turn $t$, the environment provides a global team reward $r^{\mathrm{team}}$ and an agent-specific local reward $r^{\mathrm{loc}}_{i}$ that evaluates its subtask performance. These components are combined using a hyperparameter $\alpha$ to form the final reward for agent $i$:
$r_{t,i}\;=\;\alpha\,r^{\mathrm{team}}_{t}\;+\;r^{\mathrm{loc}}_{t,i}$ (3)
This formulation balances a shared team objective with role-specific incentives. For instance, in a coder-tester MAS, the team reward $r^{\mathrm{team}}$ is the pass rate of the generated program on a set of golden unit tests. The local rewards $r^{\mathrm{loc}}_{i}$ are tailored to each role: the coder is rewarded for its own code’s pass rate, while the tester is rewarded based on the pass rate of a golden reference implementation against its generated tests. Detailed reward designs for all tasks are provided in Appendix C.1.